a new mill costs $5000 and has an annual maintenance fee of $200. salvage value after 5 years is $1000. if the interest rate is 8% the present worth is nearly?
Answer #1

Initial Cost = $5000

Annual Fees =$200

Salvage Value = $1000

Life=5 years

Discount Rate=8%

PV of Salvage Value = 1000/(1+8%)^5 = 1000/1.08^5 =1000/1.4693=$680.58

PV of Annual Fees =A*(1-(1+r)^-n)/r

=200*(1-(1+8%)^-50/8%

=200*(1-1.08^-5)/0.08

=200*(1-0.6806)/0.08

=200*0.3194/0.08

=798.54

Hence Present Worth = Initial Cost+PV of annual Fee-salvage value

=5000+798.54-680.58 = $5117.96